Heather and marika recently worked at the community service day at Harvey Scott Elementary School. If it took them 3 hours working together to put bark dust down on the side of the school, and it would have taken Heather 5 hours to do it herslef, how long would it take Marika to complete the job by herself?
Answer by stanbon(75887)

Heather and marika recently worked at the community service day at Harvey Scott Elementary School. If it took them 3 hours working together to put bark dust down on the side of the school, and it would have taken Heather 5 hours to do it herslef, how long would it take Marika to complete the job by herself?
----------------------------------------------
Together DATA:
time = 3 hr/job ; rate = 1/3 job/hr.
--------------------------------------
Heather DATA:
time = 5 hr/job ; rate = 1/5 job/hr
--------------------------------------
Marika DATA:
time = x hr/job ;rate = 1/x job/hr
----------------------------------------
Equation:
rate + rate = together rate
1/5 + 1/x = 1/3
Multiply thru by 15x to get:
3x + 15 = 5x
2x = 15
x = 7 1/2 hrs (time for Marika to do the job alone)
===================
Cheers,
Stan H.